EVALUATE THE FUNCTION WHEN X=2,X=0,X=-2.
g(x)=x+4

Found 2 solutions by Nate, checkley71:
Answer by Nate(3500)

g(-2) = -2 + 4 = 2
g(0) = 0 + 4 = 4
g(2) = 2 + 4 = 6

Answer by checkley71(8403)

G(X)=X+4
G(2)=2+4
G(2)=6 SOLUTION
G(0)=0+4
G(0)=4 SOLUTION
G(-2)=-2+4
G(-2)=+2 SOLUTION